OpenStudy (anonymous):

If AB = 4 feet, BC = 17 feet, and FG = 24 feet, what is the length of EF? 5 11/17 about correct

OpenStudy (anonymous):

do u have a diagram for this qus...?

OpenStudy (anonymous):

no there's no diagram/drawing

OpenStudy (anonymous):

5 11/17

OpenStudy (anonymous):

do u know how to get this...?

OpenStudy (anonymous):

yes 4 x 24 =95/17 =5.65 i just turned it into fraction

OpenStudy (anonymous):

yeah 4/17=EF/24

OpenStudy (anonymous):

im correct then right ? thx

OpenStudy (anonymous):

welcome...!!